### v1

# A Novel Compact LLM Framework for Local, High-Privacy EHR Data Applications

Abstract
Large Language Models (LLMs) have shown impressive capabilities in natural language processing, yet their use in sensitive domains like healthcare, particularly with Electronic Health Records (EHR), faces significant challenges due to privacy concerns and limited computational resources. This paper presents a compact LLM framework designed for local deployment in settings with strict privacy requirements and limited access to high-performance GPUs. We introduce a novel preprocessing technique that uses information extraction methods, e.g., regular expressions, to filter and emphasize critical information in clinical notes, enhancing the performance of smaller LLMs on EHR data. Our framework is evaluated using zero-shot and few-shot learning paradigms on both private and publicly available (MIMIC-IV) datasets, and we also compare its performance with fine-tuned LLMs on the MIMIC-IV dataset. The results demonstrate that our preprocessing approach significantly boosts the prediction accuracy of smaller LLMs, making them suitable for high-privacy, resource-constrained applications. This study offers valuable insights into optimizing LLM performance for sensitive, data-intensive tasks while addressing computational and privacy limitations.
Keywords LLM  $\cdot$ EHR  $\cdot$ Privacy
1 Introduction
Large Language Models (LLMs), such as GPT [1] and BERT [2], have gained substantial popularity and usefulness in recent years due to their remarkable ability to understand and generate human-like text. These models leverage vast amounts of data and complex neural network architectures to deliver insights and automation across various fields. The effectiveness of LLMs in processing and producing text has made them invaluable tools in both academic research and industry applications, pushing the boundaries of what machines can achieve in natural language understanding and generation.
Electronic Health Records (EHR) are crucial for modern healthcare as they contain comprehensive patient data [3], including both structured data (e.g., demographics, medications, lab results, genomics) and unstructured data (e.g., clinical notes). The combination enhances personalized clinical decisions, drug discovery, and informative policy-making. Despite the advance in EHR systems, the lack of precise disease phenotyping remains a significant barrier to their full application. While the Internation Classification of Diseases (ICD) codes are recorded in most EHRs, they often fail to capture the details of a patient’s condition [4]. In contrast, clinical notes can more accurately reflect disease progression, but their unstructured nature makes it challenging to extract relevant information. Manual annotation for these notes is labor-intensive, prone to errors [5], and requires highly-skilled healthcare professionals, further complicating the problem. Additionally, the sensitive nature of medical records demands strict adherence to privacy regulations [6], limiting the use of cloud-based LLMs such as ChatGPT.
To tackle the complexities of highly private EHR data, LLMs present a compelling solution to automate annotation processes, reducing the labor and time required to extract disease phenotype. However, the deployment of large-scale LLMs, such as LLaMA-70b, typically requires cloud-based infrastructure, raising privacy concerns due to sensitive health data. In addition, these models demand significant computational power, particularly high-performance GPUs with large amounts of VRAM, which may not be accessible to many labs.
Given these constraints, we focus on smaller, more accessible, locally-deployed LLMs such as Gemma-7B and LLaMA-7B. Although these models reduce the computational burden, they introduce new challenges due to their smaller context window [7, 8]: Clinical notes are often lengthy and unstructured, making it difficult for these smaller models to capture relevant information effectively.
To address these challenges, we propose a novel framework that enhances the performance of smaller LLMs by incorporating a preprocessing step. This step filters and focuses on disease-relevant contexts, reducing the input length and allowing the LLMs to work within their context window limitations. By focusing on the most pertinent sections of the clinical notes, we aim to improve the accuracy of disease phenotyping, while maintaining computational feasibility and data privacy. We present a comparison between our framework and others in Figure 1.
Figure 1: A diagram representing existing methods and our method to deal with EHR dataset.
To evaluate the effectiveness of the proposed method, we apply it to extract disease phenotypes from clinical notes using both zero-shot and few-shot learning paradigms, comparing performance with and without preprocessing. We use metastasis as an example, a critical condition that is not fully captured by ICD codes but is highly important in practice. Additionally, we fine-tune existing LLMs to further assess the impact of our approach. The advantages of our preprocessing method are demonstrated through experimental analysis on two datasets: a private dataset comprising 7,284 head-and-neck cancer (HNC) patients and the publicly available Medical Information Mart for Intensive Care (MIMIC)-IV dataset [9]. Our findings show that preprocessing significantly improves the accuracy of predictions made by smaller LLMs. In summary, the contributions of our work are as follows:
  1. 1.
    Investigation of LLM Performance for EHR Data: We systematically investigate the performance of relatively small LLMs for EHR datasets, identify the challenges they face, and propose solutions to enhance their effectiveness.
  2. 2.
    Introduction of a Preprocessing Framework: We propose a preprocessing framework to filter and highlight key information in long, unstructured clinical notes before feeding into LLMs, enabling smaller LLMs to make more accurate predictions. In this paper, we use regular expressions (regex) as an example.
  3. 3.
    Evaluation Across Multiple Learning Paradigms: Our study includes extensive experiments that compare the effectiveness of zero-shot and few-shot learning paradigms, both with and without the application of preprocessing, on EHR data. And we also compare the results with fine-tuned LLMs. This analysis highlights the considerable influence of preprocessing on model performance, enabling researchers with constrained GPU resources to maintain high accuracy while adhering to strict privacy standards.
2 Related work
Keyword search and statistical phenotyping
Early works of computational phenotyping search keywords from EHRs. Specifically, they first extract disease, medication, or lab-related keywords by simple searching algorithms, and then derive the disease phenotype based on rules designed by domain experts [10]. Rule-based systems, such as Unified Medical Language System [11], were utilized for various phenotypes including diabetes mellitus [12] and adverse drug events [13]. However, the designed rules can fall short in inferring complicated clinical notes. To address this issue, previous studies utilized statistical learning methods to predict disease phenotype after feature engineering the searched keywords and structural data [14, 15, 16]. Although these methods improve the manually designed rules, they usually ignore the context of the searched keywords in the clinical notes and may misinterpret those keywords.
Fine-tuning pre-trained language models
To effectively learn the disease phenotype in the clinical notes, a popular approach is to fine-tune pre-trained language models, such as BERT or clinical BERT [17], using annotated clinical documents. However, fine-tuned models were mainly developed for clinical text mining and named entity extraction [18, 19, 20] and less intended for disease phenotyping, potentially due to the lack of publicly available annotated data. [21] first fine-tuned BERT to extract comprehensive collections of breast cancer-related phenotypes using clinical notes from a Chinese hospital. [22] then adopted a similar approach to develop a cancer domain-specific BERT model (CancerBERT) for breast cancer-related phenotypes extraction from the University of Minnesota Clinical Data Repository (21291 patients). [23] fine-tuned BERT and BART models to automate the generation of discharge summaries for neurology patients, aiming to reduce physician burnout by enhancing the efficiency of clinical documentation. [24] proposed NYUTron, a LLM pre-trained and fine-tuned on inpatient clinical notes of 387,144 patients at New York University to provide an automatic and comprehensive profiling of patients. Notably, all above applications require highly skilled clinicians to annotate notes from a large-scale Private HNC Dataset, which can be labor-intensive and time-consuming.
Unsupervised learning using LLMs
LLMs can understand contextual language and generate human-like text, excelling across a wide range of NLP tasks [25, 26]. Several studies have directly employed LLMs for various clinical tasks [27]. For example, [28] applied GPT-3.5 to extract social determinants and family history of 1,000 deidentified notes from a University Hospital. [29] employed both GPT-3.5 and GPT-4 to create draft responses for patient messages in electronic inboxes, demonstrating statistically significant decreases in the 4-item physician task load score derivative. [30] employed GPT-4 on private datasets from UW Medical Center Montlake, UW Medical Center Northwest, and Harborview Medical Center to aid clinicians in classifying perioperative risks. [31] introduced a two-agent LLM system named EHR-CoAgent to predict disease from EHR. The system comprises two LLM agents: a predictor agent that makes predictions and articulates reasoning processes, and a critic agent that reviews incorrect predictions and provides feedback for improvements. Remarkably, most of the studies utilized cloud-based LLMs such as OpenAI API instead of locally inferred models, which could result in some privacy issues in more sensitive EHRs.
3 Method
3.1 Proposed Framework
Our observations indicate that LLMs with fewer parameters often fail to capture essential information when processing extensive text inputs. This issue is compounded in settings where computational resources are constrained, such as laboratories equipped with GPUs with small Video Random Access Memory (VRAM), which limits the deployment of more robust models.
In response to these computational constraints and privacy concerns inherent in processing sensitive data such as EHR, we have developed a structured framework to enhance the analysis of clinical notes.
The initial step involves identifying the most relevant information within the clinical text. In this paper, we demonstrate the use of regular expressions (regex) as a practical example, though it is important to emphasize that other Natural Language Processing (NLP) techniques can also be employed. To illustrate the proposed framework, we focus on the diagnosis of ‘metastasis’ in cancer cohorts. This approach can be readily extended to other diagnoses and patient cohorts with clinical expertise. In diagnosing metastasis, we search for keywords such as ‘metastasis,’ ‘metastatic,’ and other related words and extract the surrounding context for further examination. The full list of “keywords” are listed in the Appendix. This preprocessing step streamlines data handling, reducing the need for labor-intensive manual annotation while enabling more efficient downstream analysis for LLMs. Subsequently, we design customized prompts that are specifically structured to aid the model in accurately identifying medical conditions such as metastasis. These prompts are utilized within both zero-shot and few-shot learning frameworks to refine the model’s predictive capabilities.
The introduction of this preprocessing step markedly enhances the prediction accuracy, a claim that is substantiated by our experiments with the Gemma-7b model, as detailed in this paper. This method not only optimizes the performance of LLMs with fewer parameters but also addresses key operational challenges in handling EHR data within privacy-constrained and resource-limited environments.
3.2 Zero-Shot and Few-Shot learning
To enhance the effectiveness of zero-shot learning, we craft the prompts to guide the large language model’s understanding and analysis of the task at hand. This involves structuring the prompt in a way that directly asks the model to assess whether indications of metastasis are present based on the content of a single clinical note. By doing so, we ensure that the model can make accurate assessments without prior training on specific examples, relying solely on its pre-existing knowledge and the context provided within the prompt.
In our few-shot learning methodology, we adopt a hands-on training approach by carefully selecting a balanced set of examples from the dataset. This collection typically consists of $n$ positive samples, indicating the presence of metastasis, $n$ negative samples, where metastasis is absent, and $n$ neutral samples that lack information relevant to the disease state. We term this selection as $n$-shot learning. These samples are thoughtfully integrated into the learning prompts. This tailored integration is designed to familiarize the large language model with the specific subtleties and requirements of the task at hand. By interacting with this curated set of examples, the model can recalibrate its approach based on the limited yet significant exposure to task-specific data. We list the prompts for both zero-shot and few-shot learning in the Appendix.
3.3 Zero-Shot and Few-Shot learning with preprocessing
The preprocessing of clinical notes starts with the segmentation of the text into individual sentences. This step is essential for the meticulous examination of each sentence on its own, allowing for the precise identification of specific contents within the notes. During this review phase, we rigorously scan each sentence for predefined keywords related to metastasis, including ‘metastasis,’ ‘metastatic,’ and so on, which are listed in the Appendix.
Upon detecting any of these keywords in a sentence, we not only flag that sentence for extraction but also include the sentences immediately before and after it. This approach ensures that we capture a comprehensive context surrounding mentions of metastasis, encompassing any preliminary or subsequent information crucial for an in-depth analysis. Extracting adjacent sentences helps preserve the continuity and relevance of the data, thereby enhancing the accuracy of subsequent interpretations. In this process, we employ regular expressions to identify and extract relevant sentences efficiently. However, more sophisticated NLP techniques could also be utilized to refine this extraction process further.
Once the preprocessing of the text is completed, we proceed with the analysis using the LLMs. We apply the same prompt that was used in the zero-shot and few-shot learning experiment, which is carefully designed to instruct the model to determine the presence of metastasis based on the preprocessed clinical note.
3.4 LLM fine tuning
We utilize the pre-trained LLMs, Gemma-2b and Gemma-7b, to classify sequences of text based on its extensive training on diverse datasets. Given the absence of definitive ground truth in our dataset, we carefully select 50 samples that explicitly include terms related to metastasis; these are designated as our positive samples. Similarly, we choose 50 samples that suggest the absence of metastatic conditions, which we classify as negative samples.
We then split these 100 curated samples into two sets: 80% of the samples are used for training the model, while the remaining 20% are set aside for validation. This division allows us to train the model on a robust set of data while also holding back a subset to test the model’s ability to generalize to new, unseen data.
To enhance the model’s adaptability and responsiveness to our specific task without extensively altering its pre-trained weights, we incorporated the Low-Rank Adaptation (LoRA, [32]) technique. LoRA allows us to dynamically adjust the model’s behavior by introducing trainable parameters that focus on specific parts of the model, such as attention and feed-forward layers. The trainable parameters is around 3.04% for Gemma-2b and 2.29% for Gemma-7b.
Once the model has been thoroughly trained and validated, we apply it to the entire dataset, not limiting its use to only the samples involved in the training process. We then utilize the outcomes from this comprehensive application for subsequent validation.
Due to privacy restrictions, we are not permitted to use the Private HNC Dataset for model fine-tuning. As a result, we exclusively use the MIMIC-IV dataset for fine-tuning the model.
4 Applications
4.1 Introduction of Two EHR Datasets
Private HNC dataset
This dataset comprises 7,284 patients of HNC, each having at least one ICD-9 or ICD-10 code along with a corresponding diagnosis date. Among these patients, 2,637 have at least one code beginning with ICD-10 codes C78, C79, or C80, representing potential metastasis. For all of the patients with the relevant code, collectively there are 97,703 records, of which 85,870 (approximately 87.9%) include at least one clinical note containing metastasis-related keywords within a 30-day window (15 days before and after the diagnosis date).
MIMIC-IV dataset
The MIMIC-IV is an extensive, anonymized EHR database encompassing data from over 70,000 patients admitted to Beth Israel Deaconess Medical Center between 2008 and 2019 [9]. This period covers more than a decade of intensive care unit (ICU) admissions, making it a pivotal resource for medical research. With 26 tables of detailed patient data including demographics, lab measurements, procedures, medications, ICD codes, and clinical notes, MIMIC-IV offers comprehensive access to real-world clinical data. This database adheres strictly to the Health Insurance Portability and Accountability Act (HIPAA), ensuring patient privacy while supporting a wide range of medical studies. The breadth and depth of MIMIC-IV facilitate large-scale analyses critical for advancing clinical practice and shaping health policy.
We filter for subjects with at least one clinical note containing metastasis-related keywords, resulting in a dataset comprising notes from 13,923 unique subjects across 28,390 hospital admissions. After examining the ICD codes, 8,510 (approximately 61.1%) of these subjects have ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, C80, representing metastasis of tumor. Additionally, 18,552 hospital admissions (approximately 65.3%) are associated with these ICD codes.
4.2 Evaluation metrics
“Time-sensitive” accuracy defined on the Private HNC dataset
Creating a definitive gold standard for evaluating experimental outcomes in this study presents a complex challenge, primarily due to the incomplete reflection of diseases through ICD codes alone. To address this and ensure a robust comparative analysis, we employ a methodological approach for our Private HNC Dataset centered around the use of ICD-10 codes C78, C79, and C80, which are indicative of metastatic conditions. It is crucial to acknowledge, however, that the absence of these codes does not unequivocally rule out the presence of metastasis in a patient.
For each patient diagnosed with metastasis according to the aforementioned codes, we review their clinical notes over a 20-day period, encompassing 10 days before and after the diagnosis date. This time frame is selected to capture the most relevant clinical observations related to the diagnosis. During this period, we conduct a systematic search for predefined keywords associated with metastasis in the clinical notes, utilizing both zero-shot and few-shot learning, with or without the preprocessing step.
The classification accuracy of our model is assessed based on the relative counts of notes: if the number of notes indicating metastasis exceeds those indicating its absence, the classification is considered correct. Conversely, if non-metastatic notes predominate, the classification is considered incorrect. Cases with an equal count are classified as inconclusive due to insufficient information. We denote the total number of correct, incorrect, and inconclusive classifications as $N_{c}$, $N_{i}$, and $N_{l}$, respectively. Additionally, if a review period lacks clinical notes altogether, we interpret this as an absence of relevant information rather than a misclassification by the LLM, and these instances are consequently excluded from our analysis. We use $P_{c}=\frac{N_{c}}{N_{c}+N_{i}+N_{l}}$ to denote the proportion of correct classifications out of all evaluated cases. Similarly, we use $P_{i}=\frac{N_{i}}{N_{c}+N_{i}+N_{l}}$ and $P_{l}=\frac{N_{l}}{N_{c}+N_{i}+N_{l}}$ to denote the proportions of incorrect and inconclusive classifications, respectively.
This methodology is designed to mitigate the inherent limitations of relying solely on ICD codes and to provide a more nuanced understanding of the LLM’s performance in classifying complex medical conditions such as metastasis. To ensure the robustness of our analysis, we extend this approach by also reviewing clinical notes over 30-day and 40-day periods, applying the same systematic evaluation for each time frame. This allows us to assess the consistency and reliability of the model’s performance across varying observation windows.
“So-called” sensitivity and specificity defined on MIMIC-IV dataset
For the MIMIC-IV dataset, the absence of timestamps for specific ICD codes precludes us from employing the same evaluative approach as with the Private HNC Dataset. Consequently, we adapt our methodology to assess the incidence of metastasis among patients through alternative means. Our modified process begins with the exclusion of all clinical notes lacking keywords pertinent to metastasis. Subsequently, we deploy both zero-shot and few-shot LLMs, applying these with and without our preprocessing techniques, to analyze the remaining notes. We also use fine-tuned LLM to classify the clinical notes.
For each patient associated with ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, or C80, we conduct a detailed analysis by tallying instances where the LLMs have classified clinical notes as indicative of metastasis. A patient is considered correctly classified if the number of notes flagged as metastatic exceeds those identified as non-metastatic. This method provides a focused evaluation of metastatic conditions within the dataset. To simplify the notation, we let $M_{LLM}^{i}$ and $N_{LLM}^{i}$ denote the group of patients classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, we let $M_{ICD}^{i}$ and $N_{ICD}^{i}$ represent the group of patients with and without metastasis based on the ICD codes, respectively. We then calculate #$\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$ to denote the proportion of patients correctly identified as having metastasis by the LLM out of all patients with a metastatic diagnosis according to ICD codes, where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality (number of elements) of a set.
We extend this approach to individual hospital admissions, allowing for a more granular assessment of the LLMs’ performance across different patient encounters. We use $M_{LLM}^{ha}$ and $N_{LLM}^{ha}$ to denote the group of hospital admissions classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, $M_{ICD}^{ha}$ and $N_{ICD}^{ha}$ represent the group of admissions with and without metastasis based on the ICD codes, respectively. We then calculate #$\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$ to denote the proportion of hospital admissions correctly identified as having metastasis by the LLM out of all admissions diagnosed with metastasis according to ICD codes.
The analysis above serves as a measure of “so-called” sensitivity. It is crucial to recognize that even an ideal model is unlikely to achieve 100% sensitivity, as clinicians may not always explicitly document metastasis in clinical notes, despite assigning a corresponding ICD code. Nevertheless, the observed trends offer valuable insights into the model’s performance and its effectiveness in these complex scenarios. Conversely, the absence of an ICD code does not necessarily indicate that metastasis is not mentioned in the notes. Ideally, for patients lacking the relevant ICD codes, the LLMs should more frequently classify them as either not having metastasis or lacking sufficient information. To assess this, we calculate the “so-called” specificity as the proportion of patients without the relevant ICD codes who are correctly classified by the LLMs as non-metastatic or inconclusive, represented by #$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$ for individual patients and #$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$ for hospital admissions.
4.3 Results
Time range Method Preprocessing status Classified correctly ($P_{c}$) Classified incorrectly ($P_{i}$) No information is provided ($P_{l}$)
20 days Zero-shot Preprocessed 0.8796 0.0417 0.0788
No preprocessing 0.4893 0.0230 0.4878
Three-shot Preprocessed 0.8520 0.0343 0.1137
No preprocessing 0.2062 0.0138 0.7800
Six-shot Preprocessed 0.2673 0.0251 0.7077
No preprocessing 0.0714 0.0132 0.9154
30 days Zero-shot Preprocessed 0.8891 0.0443 0.0666
No preprocessing 0.5428 0.2588 0.4313
Three-shot Preprocessed 0.8697 0.0365 0.0937
No preprocessing 0.2436 0.0158 0.7406
Six-shot Preprocessed 0.3085 0.0295 0.6620
No preprocessing 0.0846 0.0154 0.9000
40 days Zero-shot Preprocessed 0.8958 0.0454 0.0587
No preprocessing 0.5792 0.0279 0.3929
Three-shot Preprocessed 0.8781 0.0375 0.0843
No preprocessing 0.2706 0.0173 0.7121
Six-shot Preprocessed 0.3368 0.0326 0.6306
No preprocessing 0.0992 0.0177 0.8832
Table 1: Classification results with different preprocessing and training conditions for our Private HNC Dataset.
Preprocessing improves classification accuracy for the private HNC dataset
Table 1 represents the experiment results using our Private HNC Dataset. The results from our investigations substantiate the efficacy of the preprocessing steps integrated into our framework. The data clearly indicates that this preprocessing significantly enhances the classification accuracy across various modeling paradigms, encompassing both zero-shot and few-shot learning approaches. This enhancement is particularly crucial when employing compact, locally-operated LLMs for processing EHR datasets. Such models, due to their smaller size relative to their more extensive counterparts, often face computational and memory constraints that can limit their effectiveness.
Furthermore, our analysis reveals a noteworthy trend concerning few-shot learning: as the number of examples included in the training increases, there is a discernible degradation in model performance. This phenomenon can be attributed to the input length sensitivity of smaller-scale LLMs. Specifically, as the input length increases with the addition of more examples, these models tend to lose focus on extracting and processing the key information critical for accurate classification. This issue underscores the challenges posed by the inherent limitations of smaller LLMs when tasked with processing complex and lengthy datasets such as EHRs.
By mitigating these challenges through our targeted preprocessing steps, we are able to considerably improve the utility of smaller LLMs in high-privacy, local computational environments, thereby enhancing their applicability and effectiveness in real-world clinical settings. This finding not only advances our understanding of model scalability and efficiency but also opens avenues for further research into optimizing LLM frameworks for specialized data-intensive tasks in constrained environments.
Preprocessing enhances sensitivity while maintaining high specificity on the MIMIC-IV dataset
Method Preprocessing status
Results (Patient)
Results (Hospital admission)
Zero-shot Preprocessed 0.7852 0.7237
No preprocessing 0.0511 0.0229
Three-shot Preprocessed 0.6223 0.4801
No preprocessing 0.0149 0.0063
Six-shot Preprocessed 0.4404 0.2799
No preprocessing 0.0043 0.0016
Classification - 2b Preprocessed 0.2972 0.3710
No preprocessing 0.2754 0.2387
Classification - 7b Preprocessed 0.1816 0.2657
No preprocessing 0.0009 0.0011
Table 2: LLM classification results for “so-called” sensitivity analysis for MIMIC-IV dataset with different preprocessing and shot settings. Results (Patient) denotes Percentage of patients with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$). The rows of 2b and 7b are for Gemma-2b and for Gemma-7b.
The results in Table 2 highlight the substantial impact of preprocessing on the performance of LLMs when applied to clinical notes from the MIMIC-IV dataset across both zero-shot and few-shot learning settings.
Preprocessing leads to significant improvements in classification accuracy across various learning settings. In zero-shot learning on a per-patient basis, preprocessing results in a substantial increase in accuracy compared to models without preprocessing. Similarly, both few-shot settings show large performance gains when preprocessing is applied. This pattern holds true in evaluations based on hospital admissions, where the preprocessed zero-shot model performs far better than the model without preprocessing. The same trend is observed in few-shot models, with preprocessing consistently yielding significantly higher accuracy than models that do not use preprocessing.
Another notable trend, similar to our findings with the private dataset, is that the model’s performance tends to decline as the number of provided examples increases. And it is likely because longer prompts with more examples make it difficult for relatively small models to focus on the key information in the clinical notes.
We also observe that with a limited number of samples, the combination of preprocessing and zero-shot learning outperforms fine-tuning the model on those samples before classification. This suggests that preprocessing can be more effective in guiding the model to relevant information than relying on a small amount of fine-tuning data.
Table 2 focuses solely on the “so-called” sensitivity (i.e., power) of the method. Additional Table 3 in the Appendix provide further validation by assessing the “so-called” specificity. This table shows that all methods perform well in terms of specificity. As a result, the primary emphasis should be on sensitivity, where the preprocessed dataset outperforms the others.
Summary
All methods demonstrate high specificity and accurately identify non-metastatic cases, the key advantage lies in the models’ sensitivity. The overall conclusion from these results is that preprocessing, which extracts and highlights critical information from clinical notes, significantly improves LLM performance. This makes preprocessing an essential step when using smaller LLMs on unstructured and complex datasets like EHRs. However, as the number of examples (shots) increases, the performance tends to decline, probably due to input length limitations in smaller models. This highlights the importance of preprocessing for maintaining model efficiency in resource-constrained environments. Furthermore, the findings suggest that when ground truth data is scarce, combining preprocessing with zero-shot learning may be more effective than fine-tuning the model with a limited number of samples. This indicates that in situations where annotation is labor-intensive, particularly for EHR datasets, using preprocessing with zero-shot learning is a more efficient and accurate approach than manual annotation.
5 Discussion
In this study, we introduce a novel compact LLM framework tailored for local deployment on EHR data, addressing significant challenges such as privacy concerns and computational limitations in healthcare settings. Our approach, which integrates a preprocessing step such as regular expressions, significantly enhances the performance of compact LLMs. This preprocessing not only reduces the data’s complexity before it reaches the LLM but also ensures that the models focus on the most relevant information, thereby improving accuracy and efficiency.
Using metastasis as an example, we demonstrate the importance of preprocessing using the Private HNC dataset and MIMIC-IV dataset. The results of our experiments underscore the potential of compact LLMs to perform robustly even under constraints, aligning with previous research that emphasizes the scalability and adaptability of LLMs in various domains. However, unlike previous studies that primarily focus on maximizing model performance without regard for computational efficiency, our work specifically targets the balance between performance and resource utilization, which is critical in many practical healthcare applications.
Our study offers practical insights for the local deployment of LLMs, particularly when utilizing models of a smaller scale. We recommend prioritizing the preprocessing of datasets to enhance model performance rather than focusing primarily on fine-tuning. This approach facilitates more effective use of limited computational resources and maintains data privacy, proving especially advantageous in settings where extensive model training is impractical.
While our findings are promising, the study is not without limitations. The framework’s reliance on regex-based preprocessing may not capture all nuances of clinical language, potentially overlooking subtle but clinically relevant information not explicitly highlighted by the regex patterns. Additionally, the performance of our framework is still dependent on the quality and scope of the input data, which may vary across different healthcare settings.
Future research could explore the integration of more sophisticated NLP techniques, such as context-aware embeddings or deep learning-based entity recognition, to enhance the accuracy and adaptability of the preprocessing step. Further studies could also evaluate the framework’s performance across a broader range of diseases and clinical conditions to better understand its generalizability.
Appendix A Additional table for specificity analysis
Method Processing status
Results (Patient)
Results (Hospital admission)
Zero-shot Preprocessed 0.9892 0.9889
No preprocessing 0.9997 0.9998
Three-shot Preprocessed 0.9933 0.9937
No preprocessing 0.9998 0.9999
Six-shot Preprocessed 0.9957 0.9965
No preprocessing 0.9999 0.9999
Classification - 2b Preprocessed 0.9732 0.9793
No preprocessing 0.9748 0.9781
Classification - 7b Preprocessed 0.9691 0.9753
No preprocessing 0.9637 0.9694
Table 3: LLM classification results for “so-called” specificity analysis for MIMIC-IV dataset with different preprocessing and shot settings. Results (Patient) denotes Percentage of patients without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$).
Appendix B Implementation details
The experimental system is equipped with an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. For zero-shot and few-shot learning, we utilize the Gemma-7b-it model, while fine-tuning is performed using both Gemma-2b-it and Gemma-7b-it models.
Predictions using zero-shot or few-shot paradigms typically take 1 to 2 weeks due to the sequential processing of individual records. Fine-tuning the models on 100 samples is significantly faster, requiring less than 10 minutes, and subsequent classification of the entire dataset takes approximately 2 days.
The implementation code for the MIMIC-IV dataset is available publicly on https://anonymous.4open.science/r/LLM_EHR-175A.
Appendix C Keywords Related to Metastasis
This section lists the keywords used to identify mentions of metastasis in clinical notes. These terms represent various aspects or stages of metastatic spread, and are essential for filtering and extracting relevant textual content during preprocessing:
  • Metastasis and Metastatic: Primary terms used to describe the spread of cancer cells from the original site to other parts of the body.
  • Metastasize, Metastases, and Metastasized: Actions and conditions involving the process of metastasis.
  • Dissemination and Distant spread: Terms that highlight the extent of cancer spread beyond local regions.
  • Metachronous, Hematogenous spread, and Lymphatic spread: Specific pathways and timings of metastatic spread.
  • Micrometastases and Infiltration: Refer to tiny, often microscopic spread of cancer cells and their penetration into surrounding tissues.
  • Tumor spread and Extra-nodal extension: Describe the progression of the tumor and the extension beyond the lymph nodes.
Appendix D Prompt used for prediction
D.1 Zero-shot learning
Zero-shot learning task Task: Classify the presence of metastasis from patient clinical notes. Respond only with the following numerical codes based on the notes provided: - (1) Yes: The notes explicitly confirm the patient has metastasis. - (2) No: The notes explicitly confirm the patient does not have metastasis. - (3) Unknown: The notes do not contain sufficient information to determine the patient’s metastasis status. Instructions: 1. Do not provide explanations or reasons for your classification. 2. Use only the information in the notes for your classification. 3. If the notes are ambiguous or lack details regarding metastasis, choose "(3) Unknown". 4. Only provide a single numerical code as the response. Examples: - Example 1: "The CT scan shows multiple nodules in the liver consistent with metastasis." - (1) - Example 2: "Patient has a history of cancer but no evidence of metastatic disease on recent imaging." - (2) - Example 3: "Patient treated for localized breast absent of metastasis." - (3) Analyze the following clinical notes to determine if the patient has metastasis: {INSERT CURRENT CLINICAL NOTES HERE}
D.2 Few-shot learning
Few-shot learning task Task: Classify the presence of metastasis from patient clinical notes. Respond only with the following numerical codes based on the notes provided: - (1) Yes: The notes explicitly confirm the patient has metastasis. - (2) No: The notes explicitly confirm the patient does not have metastasis. - (3) Unknown: The notes do not contain sufficient information to determine the patient’s metastasis status. Instructions: 1. Do not provide explanations or reasons for your classification. 2. Use only the information in the notes for your classification. 3. If the notes are ambiguous or lack details regarding metastasis, choose "(3) Unknown". 4. Only provide a single numerical code as the response. Examples: {INSERT THE EXAMPLES HERE} Analyze the following clinical notes to determine if the patient has metastasis: {INSERT CURRENT CLINICAL NOTES HERE}