### v2

# Enhancing LLMs with Smart Preprocessing for EHR Analysis

Abstract
Large Language Models (LLMs) have demonstrated remarkable proficiency in natural language processing; however, their application in sensitive domains such as healthcare, especially in processing Electronic Health Records (EHRs), is constrained by limited computational resources and privacy concerns. This paper introduces a compact LLM framework optimized for local deployment in environments with stringent privacy requirements and restricted access to high-performance GPUs. Our approach leverages simple yet powerful preprocessing techniques, including regular expressions (regex) and Retrieval-Augmented Generation (RAG), to extract and highlight critical information from clinical notes. By pre-filtering long, unstructured text, we enhance the performance of smaller LLMs on EHR-related tasks. Our framework is evaluated using zero-shot and few-shot learning paradigms on both private and publicly available datasets (MIMIC-IV), with additional comparisons against fine-tuned LLMs on MIMIC-IV. Experimental results demonstrate that our preprocessing strategy significantly supercharges the performance of smaller LLMs, making them well-suited for privacy-sensitive and resource-constrained applications. This study offers valuable insights into optimizing LLM performance for local, secure, and efficient healthcare applications. It provides practical guidance for real-world deployment for LLMs while tackling challenges related to privacy, computational feasibility, and clinical applicability.
Keywords Large Language Models, Privacy-preserving Machine Learning, Data Mining, Local Processing, Electronic Health Records, Regular Expression, Retrieval-Augmented Generation
1 Introduction
Large Language Models (LLMs), such as GPT [1] and BERT [2], have gained substantial popularity and importance in recent years due to their remarkable ability to understand and generate human-like text. These models leverage vast amounts of data and complex neural network architectures to deliver insights and automation across various fields. The effectiveness of LLMs in processing and producing text has made them invaluable tools in both academic research and industry applications, pushing the boundaries of what machines can achieve in natural language understanding and generation.
Electronic Health Records (EHR) are crucial for modern healthcare as they contain comprehensive patient data [3], including both structured data (e.g., demographics, medications, lab results, genomics) and unstructured data (e.g., clinical notes). The combination enhances personalized clinical decisions, drug discovery, and informative policy-making. Despite the advances in EHR systems, the lack of precise disease phenotyping remains a major challenge for their full utilization. While the International Classification of Diseases (ICD) codes are recorded in most EHRs, they often fail to capture the details of a patient’s condition [4]. In contrast, clinical notes can more accurately reflect disease progression, but their unstructured nature makes it challenging to extract relevant information. Manual annotation for these notes is labor-intensive, prone to errors [5], and requires highly skilled healthcare professionals, further complicating the problem. Additionally, the sensitive nature of medical records demands strict adherence to privacy regulations [6], limiting the use of cloud-based LLMs such as ChatGPT.
To tackle the complexities of the clinical notes in highly private EHR data, LLMs present a compelling solution to automate annotation processes, reducing the labor and time required to extract disease phenotypes. However, the deployment of large-scale LLMs, such as LLaMA-70B, typically requires cloud-based infrastructure, raising privacy concerns due to sensitive health data. In addition, these models demand significant computational power, particularly high-performance GPUs with large amounts of Video Random Access Memory (VRAM), which may not be accessible to many labs.
Given these constraints, we focus on smaller, more accessible, locally-deployed LLMs such as Gemma-7B and LLaMA-7B. Although these models reduce the computational burden, they introduce new challenges due to their smaller context window [7, 8]: Clinical notes are often lengthy and unstructured, making it difficult for these smaller models to capture relevant information effectively.
To address these challenges, we propose a novel framework that enhances the performance of smaller LLMs by incorporating a preprocessing step. This step filters and focuses on disease-relevant contexts, reducing the input length and allowing the LLMs to work within their context window limitations. By focusing on the most pertinent sections of the clinical notes, we aim to improve the accuracy of disease phenotyping, while maintaining computational feasibility and data privacy. We present a comparison between our framework and others in Figure 1. Compared to other methods, our framework incorporates a specifically designed preprocessing step to extract the most relevant information, demonstrated in this paper through regular expressions (regex), which identify predefined relevant keywords, and Retrieval-Augmented Generation (RAG), which retrieves sentences based on semantic similarity when prior knowledge is limited. We should note that these are just two examples to illustrate the power of preprocessing in our framework, and other NLP methods can be easily integrated into the framework.
Figure 1: A diagram representing existing methods and our method to deal with the EHR dataset.
To evaluate the effectiveness of the proposed method, we apply it to extract disease phenotypes from clinical notes using both zero-shot and few-shot learning paradigms, comparing their performance with and without preprocessing. Additionally, we fine-tune existing LLMs to further evaluate the effectiveness of our approach. In this paper, we use “metastasis” as a case study—a critical condition that is not fully captured by ICD codes but holds significant clinical importance. Importantly, our framework can be easily extended to other medical conditions and terminologies, making it broadly applicable across various healthcare domains. The advantages of our preprocessing method are demonstrated through experimental analysis on two datasets: a private dataset comprising 7,284 head-and-neck cancer (HNC) patients and the publicly available Medical Information Mart for Intensive Care (MIMIC)-IV dataset [9]. To further assess stability, we repeat the evaluation on a subset of the MIMIC-IV dataset ten times. Our findings show that preprocessing significantly improves the accuracy of predictions made by smaller LLMs. In summary, the contributions of our work are as follows:
  1. 1.
    Investigation of LLM Performance for EHR Data: We systematically investigate the performance of relatively small LLMs for EHR datasets, identify the challenges they face, and propose solutions to enhance their effectiveness.
  2. 2.
    Introduction of a Preprocessing Framework: We propose a preprocessing framework to filter and highlight key information in long, unstructured clinical notes before feeding into LLMs, enabling smaller LLMs to make more accurate predictions. By leveraging simple yet effective tricks like regex-based filtering and RAG, we demonstrate significant performance improvements in resource-constrained settings. To the best of our knowledge, no previous work has highlighted the effectiveness of simple regex-based approaches in this context.
  3. 3.
    Evaluation Across Multiple Learning Paradigms: Our study includes extensive experiments that compare the effectiveness of zero-shot and few-shot learning paradigms, both with and without the application of preprocessing, on EHR data. Different LLMs are also evaluated to ensure robustness. We also compare the results of our framework with fine-tuned LLMs. Experimental results indicate that, in scenarios with limited annotated labels, preprocessing proves to be significantly more efficient and cost-effective than fine-tuning. This analysis highlights the considerable influence of preprocessing on model performance, enabling researchers with constrained GPU resources to maintain high accuracy while adhering to strict privacy standards.
2 Related Work
Keyword search and statistical phenotyping
Early works of computational phenotyping search keywords from EHRs. Specifically, they first extract disease, medication, or lab-related keywords by simple searching algorithms, and then derive the disease phenotype based on rules designed by domain experts [10]. Rule-based systems, such as The Unified Medical Language System [11], were utilized for various phenotypes including diabetes mellitus [12] and adverse drug events [13]. However, the designed rules can fall short in inferring complicated clinical notes. To address this issue, previous studies utilized statistical learning methods to predict disease phenotype after feature engineering the searched keywords and structural data [14, 15, 16]. Although these methods improve the manually designed rules, they usually ignore the context of the searched keywords in the clinical notes and may misinterpret those keywords.
Fine-tuning pre-trained language models
To effectively learn the disease phenotype in the clinical notes, a popular approach is to fine-tune pre-trained language models, such as BERT or clinical BERT [17], using annotated clinical documents. However, fine-tuned models were mainly developed for clinical text mining and named entity extraction [18, 19, 20] and less intended for disease phenotyping, potentially due to the lack of publicly available annotated data. [21] first fine-tuned BERT to extract comprehensive collections of breast cancer-related phenotypes using clinical notes from a Chinese hospital. [22] then adopted a similar approach to developing a cancer domain-specific BERT model (CancerBERT) for breast cancer-related phenotypes extraction from the University of Minnesota Clinical Data Repository (21291 patients). [23] fine-tuned BERT and BART models to automate the generation of discharge summaries for neurology patients, aiming to reduce physician burnout by enhancing the efficiency of clinical documentation. [24] proposed NYUTron, an LLM pre-trained and fine-tuned on inpatient clinical notes of 387,144 patients at New York University to provide an automatic and comprehensive profiling of patients. Notably, all above applications require highly skilled clinicians to annotate notes from a large-scale Private HNC Dataset, which can be labor-intensive and time-consuming.
Unsupervised learning using LLMs
LLMs can understand contextual language and generate human-like text, excelling across a wide range of NLP tasks [25, 26]. Several studies have directly employed LLMs for various clinical tasks [27]. For example, [28] applied GPT-3.5 to extract social determinants and family history of 1,000 deidentified notes from a university hospital. [29] employed both GPT-3.5 and GPT-4 to generate draft responses for patient messages in electronic inboxes, demonstrating statistically significant decreases in the 4-item physician task load score derivative. [30] employed GPT-4 on private datasets from UW Medical Center Montlake, UW Medical Center Northwest, and Harborview Medical Center to assist clinicians in classifying perioperative risks. [31] introduced a two-agent LLM system named EHR-CoAgent to predict disease from EHR data. The system comprises two LLM agents: a predictor agent that makes predictions and articulates reasoning processes, and a critic agent that reviews incorrect predictions and provides feedback for improvements. Remarkably, most of the studies utilized cloud-based LLMs such as OpenAI API rather than locally inferred models, which could result in some privacy issues in more sensitive EHRs.
3 Methods
3.1 Proposed Framework
Our observations indicate that LLMs with fewer parameters often fail to capture essential information when processing extensive text inputs. This issue is compounded in settings where computational resources are constrained, such as laboratories equipped with GPUs with small VRAM, which limits the deployment of more robust models.
In response to these computational constraints and privacy concerns inherent in processing sensitive data such as EHR, we have developed a structured framework to enhance the analysis of clinical notes. The initial step involves identifying the most relevant information within the clinical text. In this paper, we demonstrate the use of regex and RAG as practical examples, though it is important to emphasize that other Natural Language Processing (NLP) techniques can also be employed. To illustrate the proposed framework, we focus on the diagnosis of “metastasis” in cancer cohorts. This approach can be readily extended to other diagnoses and patient cohorts.
First, we emphasize how we use regex to do the preprocessing procedure with clinical expertise. In diagnosing metastasis, we search for keywords such as “metastasis,” “metastatic,” and other related words and extract the surrounding context for further examination. The full list of “keywords” are listed in the Appendix. This preprocessing step streamlines data handling, reducing the need for labor-intensive manual annotation while enabling more efficient downstream analysis for LLMs. Additionally, when prior knowledge is insufficient—such as not knowing the associated terms for “metastasis”—we turn to the RAG approach. RAG combines LLMs with information retrieval systems, retrieving sentences that are semantically similar to input keywords [32] based on cosine similarities. The introduction of these preprocessing steps markedly enhances prediction accuracy, a claim substantiated by our experiments with the Gemma-7B model and other LLMs fine-tuned with biomedical knowledge, as detailed in this paper.
After the preprocessing step, we design customized prompts that are specifically structured to aid the model in accurately identifying medical conditions such as metastasis. These prompts are utilized within both zero-shot and few-shot learning frameworks to refine the model’s predictive capabilities.
Figure 2: Flowchart of our proposed framework and evaluation process for two EHR datasets.
3.2 Evaluations with Zero-Shot and Few-Shot Learning
To evaluate the effectiveness of zero-shot learning, we craft the prompts to guide the large language model’s understanding and analysis of the task at hand. This involves structuring the prompt in a way that directly asks the model to assess whether indications of metastasis are present based on the content of a single clinical note. By doing so, we ensure that the model can make accurate assessments without prior training on specific examples, relying solely on its pre-existing knowledge and the context provided within the prompt.
In our few-shot learning methodology, we adopt a hands-on training approach by carefully selecting a balanced set of examples from the dataset. This collection typically consists of $n$ positive samples, indicating the presence of metastasis, $n$ negative samples, where metastasis is absent, and $n$ neutral samples that lack information relevant to the disease state. We term this selection as “$n$-shot” learning. These samples are thoughtfully integrated into the learning prompts. This integration is designed to familiarize the LLMs with the specific subtleties and requirements of the task at hand. By interacting with this curated set of examples, the model can recalibrate its approach based on the limited yet significant exposure to task-specific data. We list the prompts for both zero-shot and few-shot learning in the Appendix.
We start by processing the entire dataset once using the zero-shot and few-shot learning paradigm to establish overall performance. To further assess stability and variability, we repeat the procedure 10 times using different random seeds, testing only a subset of the MIMIC-IV data due to time constraints. Additionally, we investigate whether varying shot settings improve performance across different LLMs and evaluate output consistency under different conditions. To ensure a comprehensive comparison, we experiment with multiple LLMs, including Gemma-7B, as well as medically fine-tuned models such as Bio-Medical-LLaMA-3-8B and LLaMA-2-7B-Chat-Med.
3.3 Evaluations with Zero-Shot and Few-Shot Learning with Preprocessing
The preprocessing of clinical notes starts with the segmentation of the text into individual sentences. This step is essential for the meticulous examination of each sentence on its own, allowing for the precise identification of specific contents within the notes. After that, we use regex and RAG as two examples to extract the relevant information. Under the regex-based approach, we systematically scan each sentence for predefined keywords related to metastasis, including “metastasis,” “metastatic” and so on, which are listed in the Appendix. Upon detecting any of these keywords in a sentence, we not only flag that sentence for extraction but also include the sentences immediately before and after it. This approach ensures that we capture a comprehensive context surrounding mentions of metastasis, encompassing any preliminary or subsequent information crucial for an in-depth analysis. Extracting adjacent sentences helps preserve the continuity and relevance of the data, thereby enhancing the accuracy of subsequent interpretations. In addition to regex, we utilize the RAG approach when prior knowledge is insufficient or when defining keywords requires domain expertise. Instead of manually determining which sentences are relevant to metastasis, RAG selects and retrieves sentences solely based on similarity scores, measured by Cosine Similarity [33]. This approach provides a viable solution in scenarios where predefined keyword lists are not available.
Once the preprocessing of the text is completed, we investigate whether varying processing settings improve performance on different LLMs and evaluate output consistency under different conditions. We apply the same prompt that was used in the zero-shot and few-shot learning experiment, which is carefully designed to instruct the model to determine the presence of metastasis based on the preprocessed clinical note.
For the regex-based approach, we first process the entire dataset and then conduct 10 additional runs using different random seeds on a subset of the MIMIC-IV dataset to evaluate stability and variability. Additionally, for the RAG-based approach, we repeat the procedure 10 times with different random seeds on the same subset of the MIMIC-IV dataset due to time constraints. Additionally, we assess the impact of extracting varying numbers of relevant sentences in the RAG-based approach to analyze its effect on performance. Similar to the non-preprocessed setting, both approaches are tested across multiple LLMs and different shot settings to ensure consistency and robustness across various models.
3.4 Evaluations with Fine-Tuned LLMs
To compare the effectiveness of preprocessing with the widely used fine-tuning approach, we fine-tune pre-trained LLMs, Gemma-2B and Gemma-7B, for text classification based on their extensive training across diverse datasets. Due to the lack of definitive ground truth in our dataset, we carefully curate 50 samples explicitly containing metastasis-related terms as positive samples. Likewise, we select 50 samples that indicate the absence of metastatic conditions, classifying them as negative samples. We then split these 100 curated samples into two sets: 80% of the samples are used for training the model, while the remaining 20% are set aside for validation. This division allows us to train the model on a robust set of data while also holding back a subset to test the model’s ability to generalize to new, unseen data.
To enhance the model’s adaptability and responsiveness to our specific task without extensively altering its pre-trained weights, we incorporated the Low-Rank Adaptation (LoRA, [34]) technique. LoRA allows us to dynamically adjust the model’s behavior by introducing trainable parameters that focus on specific parts of the model, such as attention and feed-forward layers. The trainable parameters are around 3.04% for Gemma-2B and 2.29% for Gemma-7B.
Once the model has been thoroughly trained and validated, we apply it to the entire dataset, not limiting its use to only the samples involved in the training process. We then utilize the outcomes from this comprehensive application for subsequent validation. Due to privacy restrictions, we are not permitted to use the Private HNC Dataset for model fine-tuning. As a result, we exclusively use the MIMIC-IV dataset for fine-tuning the model.
4 Applications
4.1 Introduction of Two EHR Datasets
Private HNC dataset
This dataset comprises 7,284 patients of HNC, each having at least one ICD-9 or ICD-10 code along with a corresponding diagnosis date. Among these patients, 2,637 have at least one code beginning with ICD-10 codes C78, C79, or C80, representing potential metastasis. For all of the patients with the relevant code, collectively there are 97,703 records, of which 85,870 (approximately 87.9%) include at least one clinical note containing metastasis-related keywords within a 30-day window (15 days before and after the diagnosis date).
MIMIC-IV dataset
The MIMIC-IV dataset is an extensive, anon-ymized EHR database encompassing data from over 70,000 patients admitted to Beth Israel Deaconess Medical Center between 2008 and 2019 [9]. This period covers more than a decade of intensive care unit (ICU) admissions, making it a pivotal resource for medical research. With 26 tables of detailed patient data including demographics, lab measurements, procedures, medications, ICD codes, and clinical notes, MIMIC-IV offers comprehensive access to real-world clinical data. This database adheres strictly to the Health Insurance Portability and Accountability Act (HIPAA), ensuring patient privacy while supporting a wide range of medical studies. The breadth and depth of MIMIC-IV facilitate large-scale analyses critical for advancing clinical practice and shaping health policy.
We filter for subjects with at least one clinical note containing metastasis-related keywords, resulting in a dataset comprising notes from 13,923 unique subjects across 28,390 hospital admissions. After examining the ICD codes, 8,510 (approximately 61.1%) of these subjects have ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, C80, representing metastasis of tumors. Additionally, 18,552 hospital admissions (approximately 65.3%) are associated with these ICD codes.
4.2 Evaluation Metrics
“Time-sensitive” sensitivity defined on the Private HNC dataset
Creating a definitive gold standard for evaluating experimental outcomes in this study presents a complex challenge, primarily due to the incomplete reflection of diseases through ICD codes alone. However, given the large volume of clinical notes, expert validation would be both labor-intensive and impractical. To address this and ensure a robust comparative analysis, we employ a methodological approach for our Private HNC Dataset centered around the use of ICD-10 codes C78, C79, and C80, which are indicative of metastatic conditions. The evaluation process is illustrated in Figure 2. Guided by expert input, these codes offer a relatively reliable means of determining a patient’s metastasis status. Therefore, we utilize them as a gold standard to evaluate the effectiveness of various methods. However, it is crucial to acknowledge that the absence of these codes does not unequivocally rule out the presence of metastasis in a patient.
For each patient diagnosed with metastasis according to the aforementioned codes, we review their clinical notes over a 20-day period, encompassing 10 days before and after the diagnosis date. This time frame is chosen to capture the most relevant clinical observations surrounding the diagnosis. Within this period, we systematically search for predefined metastasis-related keywords using regex in the clinical notes as a preprocessing step. We then apply both zero-shot and few-shot learning approaches, with and without the above preprocessing step, to assess the model’s ability to classify metastasis accurately.
The classification accuracy of our model—more precisely, its sensitivity—is assessed based on the relative counts of notes. We do not include a “time-sensitive” specificity analysis, as the absence of an ICD code related to metastasis does not guarantee that no mention of metastasis appears in the clinical notes within a specific time frame. Instead, a specificity analysis is conducted using the MIMIC-IV dataset to validate the robustness of our model. To assess sensitivity, we analyze all clinical notes within the specified time frame. If the number of notes classified by the LLM as indicating metastasis exceeds those indicating its absence, the classification is considered correct. Conversely, if non-metastatic notes predominate, the classification is deemed incorrect. Cases with an equal count are classified as inconclusive due to insufficient information. We denote the total number of correct, incorrect, and inconclusive classifications as $N_{c}$, $N_{i}$, and $N_{l}$, respectively. Additionally, if a review period lacks clinical notes altogether, we interpret this as an absence of relevant information rather than a misclassification by the LLM, and these instances are consequently excluded from our analysis. We use $P_{c}=\frac{N_{c}}{N_{c}+N_{i}+N_{l}}$ to denote the proportion of correct classifications out of all evaluated cases. Similarly, we use $P_{i}=\frac{N_{i}}{N_{c}+N_{i}+N_{l}}$ and $P_{l}=\frac{N_{l}}{N_{c}+N_{i}+N_{l}}$ to denote the proportions of incorrect and inconclusive classifications, respectively.
This methodology is designed to mitigate the inherent limitations of relying solely on ICD codes and to provide a more nuanced understanding of the LLM’s performance in classifying complex medical conditions such as metastasis. To ensure the robustness of our analysis, we extend this approach by also reviewing clinical notes over 30-day and 40-day periods, applying the same systematic evaluation for each time frame. This allows us to assess the consistency and reliability of the model’s performance across varying observation windows.
“So-called” sensitivity and specificity defined on MIMIC-IV dataset
For the MIMIC-IV dataset, the absence of timestamps for specific ICD codes prevents us from using the same evaluative approach as in the Private HNC Dataset. As a result, we evaluate the results on a per-individual or per-hospital-admission basis. The per-individual evaluation is also depicted in Figure 2. Our process begins by excluding all clinical notes that do not contain keywords related to metastasis. We then apply both regex-based and RAG-based methods to extract metastasis-related sentences for each patient, as preprocessing. Next, we deploy zero-shot and few-shot LLMs, with and without the preprocessing procedure mentioned above, to analyze the remaining notes. We also test various LLMs, such as Gemma-7B-it and the domain-specific LLaMA-2-7B-Chat-Med and Bio-Medical-LLaMA-3-8B, to broaden the evaluation. To ensure robustness, we perform ten additional runs on a subset of the dataset, evaluated exclusively on a per-individual basis. Additionally, we use fine-tuned LLMs to classify the clinical notes, allowing for a comprehensive comparison of different approaches.
We begin by outlining our evaluation method on a per-individual basis. For each patient associated with ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, or C80, we conduct a detailed analysis by tallying instances where the LLMs have classified clinical notes as indicative of metastasis. A patient is considered correctly classified if the number of notes flagged as metastatic exceeds those identified as non-metastatic. This method provides a focused evaluation of metastatic conditions within the dataset. To simplify the notation, we let $M_{LLM}^{i}$ and $N_{LLM}^{i}$ denote the group of patients classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, we let $M_{ICD}^{i}$ and $N_{ICD}^{i}$ represent the group of patients with and without metastasis based on the ICD codes, respectively. We then calculate $P_{c}^{i}=\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$ to denote the proportion of patients correctly identified as having metastasis by the LLM out of all patients with a metastatic diagnosis according to ICD codes, where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality (number of elements) of a set.
We extend this approach to individual hospital admissions, allowing for a more granular assessment of the LLMs’ performance across different patient encounters. We use $M_{LLM}^{ha}$ and $N_{LLM}^{ha}$ to denote the group of hospital admissions classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, $M_{ICD}^{ha}$ and $N_{ICD}^{ha}$ represent the group of admissions with and without metastasis based on the ICD codes, respectively. We then calculate $P_{c}^{ha}=\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$ to denote the proportion of hospital admissions correctly identified as having metastasis by the LLM out of all admissions diagnosed with metastasis according to ICD codes.
The analysis above serves as a measure of “so-called” sensitivity. It is crucial to recognize that even an ideal model is unlikely to achieve 100% sensitivity, as clinicians may not always explicitly document metastasis in clinical notes, despite assigning a corresponding ICD code. Nevertheless, the observed trends offer valuable insights into the model’s performance and its effectiveness in these complex scenarios. Conversely, the absence of an ICD code does not necessarily indicate that metastasis is not mentioned in the notes. Ideally, for patients lacking the relevant ICD codes, the LLMs should more frequently classify them as either not having metastasis or lacking sufficient information. To assess this, we calculate the “so-called” specificity as the proportion of patients without the relevant ICD codes who are correctly classified by the LLMs as non-metastatic or inconclusive, represented by #$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$ for individual patients and #$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$ for hospital admissions. We present the per-individual evaluation flowchart in Figure 2.
4.3 Results
Preprocessing improves classification accuracy for the private HNC dataset
Figure 3(a) presents a visual summary of the experimental results on our Private HNC Dataset, while the complete numerical results are detailed in Table 1. The results from our investigations substantiate the efficacy of the preprocessing steps integrated into our framework. The results clearly indicate that the regex-based preprocessing significantly enhances the classification accuracy across various modeling paradigms, encompassing both zero-shot and few-shot learning approaches. This enhancement is particularly crucial when employing compact, locally-operated LLMs for processing EHR datasets. Such models, due to their smaller size relative to their more extensive counterparts, often face computational and memory constraints that can limit their effectiveness.
Furthermore, our analysis reveals a noteworthy trend concerning few-shot learning: as the number of examples included in the training increases, there is a discernible degradation in model performance. This phenomenon can be attributed to the input length sensitivity of smaller-scale LLMs. Specifically, as the input length increases with the addition of more examples, these models tend to lose focus on extracting and processing the key information critical for accurate classification. This issue underscores the challenges posed by the inherent limitations of smaller LLMs when tasked with processing complex and lengthy datasets such as EHRs.
By mitigating these challenges through our targeted preprocessing steps, we are able to considerably improve the utility of smaller LLMs in high-privacy, local computational environments, thereby enhancing their applicability and effectiveness in real-world clinical settings. These findings not only advance our understanding of model scalability and efficiency but also open avenues for further research into optimizing LLM frameworks for specialized data-intensive tasks in constrained environments.
Preprocessing enhances sensitivity while maintaining high specificity on the MIMIC-IV dataset
(a) Classification Results for Our Private HNC Dataset.
(b) Classification Results for the MIMIC-IV Dataset.
Figure 3: Comparison of classification results for the Private HNC dataset and MIMIC-IV dataset under different preprocessing conditions and LLM settings. (a) A bar chart displaying classification results for the Private HNC dataset under different preprocessing conditions using Gemma-7B-it. The x-axis represents processing methods (Regex and No Preprocessing), while the y-axis represents the proportion of correctly classified instances. The chart includes three shot settings (Zero-shot, Three-shot, Six-shot) across different time ranges (20 days, 30 days, and 40 days). (b) A bar chart illustrating LLM classification results for the MIMIC-IV dataset under different preprocessing conditions. The x-axis represents processing methods (Regex and No Preprocessing), and the y-axis represents the proportion of correctly classified instances. The chart compares classification accuracy across various LLM configurations, including Zero-shot, Three-shot, and Six-shot learning using Gemma-7B-it, as well as two classification models fine-tuned on Gemma-2B-it and Gemma-7B-it. Each column corresponds to a specific method, while rows distinguish between Patient and Hospital Admission categories.
The results presented in Figure 3(b), with the complete numerical details available in Table 2, underscore the significant impact of regex-based preprocessing on the performance of LLMs in analyzing clinical notes from the MIMIC-IV dataset. This improvement is observed consistently across both zero-shot and few-shot learning settings. Preprocessing leads to significant improvements in classification accuracy across various learning settings. In zero-shot learning on a per-patient basis, preprocessing results in a substantial increase in accuracy compared to models without preprocessing. Similarly, both few-shot settings show large performance gains when preprocessing is applied. This pattern holds true in evaluations based on hospital admissions, where the preprocessed zero-shot model performs far better than the model without preprocessing. The same trend is observed in few-shot models, with preprocessing consistently yielding significantly higher accuracy.
Another notable trend, similar to our findings with the private dataset, is that the model’s performance tends to decline as the number of provided examples increases. This is likely because longer prompts with more examples make it difficult for relatively small models to focus on the key information in the clinical notes. We also observe that with a limited number of samples, the combination of preprocessing and zero-shot learning outperforms fine-tuning the model on those samples before classification. This suggests that preprocessing can be more effective in guiding the model to relevant information than relying on a small amount of fine-tuning data.
Repeated evaluations on subset data yielded results that aligned closely with those obtained from the full dataset. As shown in Figure 4(a), with detailed numerical outcomes provided in Table 4, the chosen preprocessing methods consistently improved model performance across all evaluated LLMs. Note that the previous results focus solely on the “so-called” sensitivity (i.e., power) of the method. Table 3 in the Appendix provides further validation by assessing the “so-called” specificity. This table shows that all methods perform well in terms of specificity. As a result, the primary emphasis should be on sensitivity, where the preprocessed dataset outperforms the others.
(a) Comparison of Different LLMs.
(b) Comparison of Different Preprocessing Approaches.
Figure 4: Comparison of LLMs and preprocessing approaches under different shot settings based on the MIMIC-IV dataset. (a) A bar chart showing the comparison of three different language models (Gemma-7B-it, LLaMA-2-7B-Chat-Med, and Bio-Medical-LLaMA-3-8B) across three shot settings: Zero-shot, Three-shot, and Six-shot. The x-axis represents preprocessing methods, including Regex and No Preprocessing. The y-axis represents Sensitivity. Each row represents a different language model, while each facet represents a different shot setting. Bars are color-coded for different preprocessing methods, and error bars indicate the range of the results. (b) A bar chart comparing different preprocessing approaches under three kinds of shot settings (Zero-shot, Three-shot, Six-shot) for Gemma-7B-it. The x-axis represents preprocessing methods, including Regex, RAG (Top 2), RAG (Top 3), RAG (Top 6), and No Preprocessing. The y-axis represents Sensitivity. Each shot setting is displayed as a separate facet, and error bars indicate the range of the results.
Regex-based preprocessing outperforms RAG on the MIMIC-IV dataset.
Figure 4(b), with detailed numerical results shown in Table 5, presents the sensitivity results comparing the regex-preprocessed, RAG-preprocessed, and no-processing approaches. Across zero-shot, three-shot, and six-shot settings, both regex and RAG exhibit strong overall performance with good consistency (as indicated by the relatively narrow ranges of sensitivity). When using RAG, retrieving the top two sentences yields the highest mean sensitivity; performance declines considerably when retrieving six sentences, likely due to additional irrelevant content that reduces accuracy. Introducing a three-shot scenario further enhances RAG’s comprehension when retrieving the top two relevant sentences, but it still falls short of regex—particularly in the zero-shot setting, where regex achieves its best performance.
Additionally, the RAG approach presents several challenges. One major issue is its reliance on a distance parameter to measure the similarity between the query and candidate sentences for ranking or filtering. Determining an optimal distance threshold is non-trivial, as it lacks a universal reference and varies depending on different keywords and text lengths. Among the tested settings, retrieving the top two sentences yields the best performance, demonstrating RAG’s ability to extract relevant information. However, it still underperforms the regex model. A possible reason is that retrieving only two sentences may omit crucial context, while longer text inputs require additional examples to help the model generate more structured outputs. Conversely, retrieving six sentences introduces excessive irrelevant content, increasing input length and leading to reduced accuracy. Moreover, as the number of retrieved sentences grows, computational overhead rises significantly, further complicating optimization. These challenges make it difficult to establish an optimal threshold for RAG-based preprocessing. In contrast, regex remains a more reliable and broadly applicable method, particularly for smaller LLMs, as it eliminates the need for similarity measures and manual tuning of retrieval parameters.
Summary
All methods demonstrate high specificity and accurately identify non-metastatic cases, the key advantage lies in the models’ sensitivity. The overall conclusion from these results is that preprocessing, which extracts and highlights critical information from clinical notes, significantly improves LLM performance. This paper presents two examples to demonstrate the effectiveness of preprocessing, while other NLP methods may also be applied to the preprocessing step. Regex, guided by predefined references, enhances model performance by extracting precise information from clinical notes. RAG, on the other hand, remains effective in the absence of a universal reference, demonstrating adaptability across various types of unstructured clinical data. This makes preprocessing an essential step when using smaller LLMs on unstructured and complex datasets like EHRs. However, as the number of examples (shots) increases, the performance tends to decline, probably due to input length limitations in smaller models. This highlights the importance of preprocessing for maintaining model efficiency in resource-constrained environments. Furthermore, the findings suggest that when ground truth data is scarce, combining preprocessing with zero-shot learning may be more effective than fine-tuning the model with a limited number of samples. This indicates that in situations where annotation is labor-intensive, particularly for EHR datasets, using preprocessing with zero-shot learning is a more efficient and accurate approach than manual annotation.
5 Discussion
In this study, we introduce a novel compact LLM framework tailored for local deployment on EHR data, addressing significant challenges such as privacy concerns and computational limitations in healthcare settings. Our approach significantly supercharges the performance of smaller LLMs by integrating smart preprocessing techniques, including but not limited to, regex and RAG. This preprocessing not only reduces the data’s complexity before it reaches the LLM but also ensures that the models focus on the most relevant information, thereby improving accuracy and efficiency. Moreover, preprocessing serves as a form of data quality control, helping to refine and standardize input data, ultimately leading to more reliable and interpretable model outputs.
Using “metastasis” as an example, we demonstrate the importance of preprocessing on the Private HNC dataset and MIMIC-IV dataset. The results of our experiments underscore the potential of compact LLMs to perform robustly even under constraints, consistent with previous research that emphasizes the scalability and adaptability of LLMs in various domains. However, unlike previous studies that primarily focus on maximizing model performance without regard for computational efficiency, our work specifically targets the balance between performance and resource utilization, which is critical in many practical healthcare applications. Furthermore, in addition to “metastasis,” we manually assessed the framework’s performance using other keywords, including “Programmed Death-Ligand 1 (PDL1),” “Human Papillomavirus (HPV),” and “Combined Positive Score (CPS).” Although our framework produces reasonably encouraging results, the lack of a definitive gold standard in ICD codes and other resources precludes us from presenting these outcomes in the paper.
Our study provides practical insights into the local deployment of LLMs, particularly for smaller-scale models. Notably, all the compact LLMs we evaluated benefit significantly from preprocessing, highlighting its effectiveness in optimizing performance. Therefore, we recommend prioritizing preprocessing techniques to enhance model accuracy rather than relying solely on fine-tuning. Furthermore, when some domain knowledge is available, simple methods like regex can be highly effective, providing a fast and efficient solution. Notably, this knowledge often does not require much domain expertise—identifying relevant synonyms is usually sufficient, a task that can be easily handled using a free AI tool like ChatGPT. In cases where domain expertise is truly necessary, our study demonstrates that RAG can be a viable alternative. However, while RAG can still yield useful results, it tends to be less effective than regex and requires more computational resources. An additional advantage we did not emphasize earlier is the impact of preprocessing on processing time. Although it introduces an extra step, preprocessing significantly reduces input length, thereby accelerating overall performance. Both regex-based and RAG-based filtering take less than 0.3 seconds per record, with subsequent LLM inference on preprocessed records requiring under 1 second. In contrast, inference on a single unprocessed record averages around 10 seconds. For a dataset exceeding 80,000 clinical notes, this translates to a processing time of 1–2 days with preprocessing, compared to 1–2 weeks without it, marking a substantial improvement in efficiency.
While our findings are promising, our study has certain limitations. The framework’s reliance on regex-based preprocessing may not fully capture the complexities of clinical language, potentially overlooking subtle yet clinically relevant information that falls outside predefined patterns. Although RAG-based methods can help address this limitation, their effectiveness can diminish without optimizing parameters, such as the number of retrieved sentences. As a potential future direction, investigating adaptive thresholding could enhance RAG’s effectiveness. Additionally, the effectiveness of our framework depends on the quality and diversity of the input data, which can vary across different healthcare settings, although we posit that preprocessing will consistently improve results compared to non-preprocessing across all scenarios. Future research could explore the integration of more advanced NLP techniques, such as context-aware embeddings or deep learning-based entity recognition, to further enhance the accuracy and adaptability of the preprocessing step. Moreover, leveraging existing biomedical databases or knowledge graphs could facilitate more advanced preprocessing, enriching the model’s understanding of clinical concepts and relationships. Finally, assessing the framework across a broader range of diseases and clinical conditions would provide deeper insights into its generalizability and robustness.
Appendix A Additional Tables
Table 1 presents the classification sensitivity results for the Private HNC dataset, comparing different preprocessing settings and shot settings. Table 2 summarizes LLM classification sensitivity performance on the MIMIC-IV dataset, evaluating the impact of preprocessing under different shot settings. Table 3 presents the specificity analysis of LLM classification results on the MIMIC-IV dataset. The table compares preprocessed and non-preprocessed data across various shot settings. Table 4 provides a detailed breakdown of LLM performance under different shot settings, specifically comparing Gemma-7B-it, and two LLaMa models (LLaMA-2-7B-Chat-Med, and Bio-Medical-LLaMA-3-8B). The table quantifies differences in classification accuracy across multiple preprocessing conditions. Table 5 compares the effectiveness of different preprocessing approaches, including regex and RAG, under varying shot settings.
Table 1: Classification results with different preprocessing and training conditions for our Private HNC Dataset.
Time range Method Preprocessing Status Classified Correctly ($P_{c}$) Classified Incorrectly ($P_{i}$) No Information ($P_{l}$)
20 days Zero-shot Preprocessed 0.8796 0.0417 0.0788
No preprocessing 0.4893 0.0230 0.4878
Three-shot Preprocessed 0.8520 0.0343 0.1137
No preprocessing 0.2062 0.0138 0.7800
Six-shot Preprocessed 0.2673 0.0251 0.7077
No preprocessing 0.0714 0.0132 0.9154
30 days Zero-shot Preprocessed 0.8891 0.0443 0.0666
No preprocessing 0.5428 0.2588 0.4313
Three-shot Preprocessed 0.8697 0.0365 0.0937
No preprocessing 0.2436 0.0158 0.7406
Six-shot Preprocessed 0.3085 0.0295 0.6620
No preprocessing 0.0846 0.0154 0.9000
40 days Zero-shot Preprocessed 0.8958 0.0454 0.0587
No preprocessing 0.5792 0.0279 0.3929
Three-shot Preprocessed 0.8781 0.0375 0.0843
No preprocessing 0.2706 0.0173 0.7121
Six-shot Preprocessed 0.3368 0.0326 0.6306
No preprocessing 0.0992 0.0177 0.8832
Table 2: LLM classification results under different preprocessing and shot settings for the MIMIC-IV dataset.
Method Preprocessing Status Results (Patient) Results (Hospital Admission)
Zero-shot Preprocessed 0.7852 0.7237
No Preprocessing 0.0511 0.0229
Three-shot Preprocessed 0.6223 0.4801
No Preprocessing 0.0149 0.0063
Six-shot Preprocessed 0.4404 0.2799
No Preprocessing 0.0043 0.0016
Classification - 2B Preprocessed 0.2972 0.3710
No Preprocessing 0.2754 0.2387
Classification - 7B Preprocessed 0.1816 0.2657
No Preprocessing 0.0009 0.0011
Notes: Results (Patient) denotes percentage of patients with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$). The rows of 2B and 7B are for Gemma-2B and for Gemma-7B.
Table 3: LLM classification results for specificity analysis on the MIMIC-IV dataset.
Method Processing Status Results (Patient) Results (Hospital Admission)
Zero-shot Preprocessed 0.9892 0.9889
No Preprocessing 0.9997 0.9998
Three-shot Preprocessed 0.9933 0.9937
No Preprocessing 0.9998 0.9999
Six-shot Preprocessed 0.9957 0.9965
No Preprocessing 0.9999 0.9999
Classification - 2B Preprocessed 0.9732 0.9793
No Preprocessing 0.9748 0.9781
Classification - 7B Preprocessed 0.9691 0.9753
No Preprocessing 0.9637 0.9694
Notes: Results (Patient) denotes Percentage of patients without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$).
Table 4: Comparison of different LLMs under various shot settings on the MIMIC-IV dataset.
Shot Setting Processing State Gemma-7B-it LLaMA-2-7B-Chat-Med Bio-Medical-LLaMA-3-8B
Zero-shot Preprocessed with regex 0.835 (0.831-0.836) 0.749 (0.729-0.763) 0.595 (0.584-0.606)
No Preprocessing 0.080 (0.078-0.081) 0.384 (0.367-0.401) 0.495 (0.482-0.516)
Three-shot Preprocessed with regex 0.813 (0.809-0.814) 0.464 (0.438-0.494) 0.712 (0.699-0.719)
No Preprocessing 0.085 (0.073-0.091) 0.217 (0.203-0.235) 0.280 (0.271-0.291)
Six-shot Preprocessed with regex 0.616 (0.609-0.621) 0.618 (0.606-0.636) 0.546 (0.538-0.555)
No Preprocessing 0.0058 (0.0024-0.0073) 0.562 (0.536-0.580) 0.427 (0.413-0.452)
Notes: Results denote the percentage of patients with relevant ICD codes classified with metastasis by LLM out of patients with relevant ICD codes. Results are based on running each model 10 times on a 5% sample of the dataset. Values are reported as mean (range) across runs.
Table 5: Comparison of preprocessing approaches under different shot settings on MIMIC-IV dataset.
Shot Setting Preprocessing Status Mean of $P_{c}^{i}$ Range of $P_{c}^{i}$
Zero-shot Preprocessed with regex 0.8493 0.8356 - 0.8630
Preprocessed with RAG (Top 2) 0.6863 0.6712 - 0.6986
Preprocessed with RAG (Top 3) 0.6534 0.6301 - 0.6575
Preprocessed with RAG (Top 6) 0.6657 0.6438 - 0.6712
No preprocessing 0.0671 0.0548 - 0.0685
Three-shot Preprocessed with regex 0.7972 0.7945 - 0.8082
Preprocessed with RAG (Top 2) 0.7383 0.5342 - 0.7671
Preprocessed with RAG (Top 3) 0.6616 0.5342 - 0.7123
Preprocessed with RAG (Top 6) 0.4494 0.4384 - 0.4658
No preprocessing 0.10 0.0548 - 0.1233
Six-shot Preprocessed with regex 0.6685 0.6575 - 0.6849
Preprocessed with RAG (Top 2) 0.6438 0.6301 - 0.6575
Preprocessed with RAG (Top 3) 0.4055 0.3973 - 0.4110
Preprocessed with RAG (Top 6) 0.1630 0.1507 - 0.1644
No preprocessing 0.0247 0.0137 - 0.0274
Notes: Results denote the percentage of patients with relevant ICD codes classified with metastasis by LLM out of patients with relevant ICD codes. Results are based on running the Gemma-7B-it model 10 times on a 1% sample of the dataset. Values are reported as mean and range across runs.
Appendix B Implementation Details
The experimental system is equipped with an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. For zero-shot and few-shot learning, we utilize the Gemma-7B-it model, while fine-tuning is performed using both Gemma-2B-it and Gemma-7B-it models.
For Gemma-7B-it model, predictions under zero-shot or few-shot paradigms typically take 1 to 2 weeks due to the sequential processing of individual records across the entire dataset. Fine-tuning the models on 100 samples is significantly faster; it requires less than 10 minutes. The subsequent classification of the entire dataset takes approximately 2 days. The regex-based approach and the RAG-based approach (with top-2 sentence retrieval) have similar processing times while retrieving the top 6 sentences using RAG increases runtime by approximately 2.5 times. LLaMA models generally take longer than Gemma models, with Bio-Medical-LLaMA-3-8B taking roughly twice as long, but they require less GPU memory and can operate within a 16GB VRAM environment.
Appendix C Keywords Related to Metastasis
This section lists the keywords used to identify mentions of metastasis in clinical notes. These terms represent various aspects or stages of metastatic spread, and are essential for filtering and extracting relevant textual content during preprocessing:
  • Metastasis and Metastatic: Primary terms used to describe the spread of cancer cells from the original site to other parts of the body.
  • Metastasize, Metastases, and Metastasized: Actions and conditions involving the process of metastasis.
  • Dissemination and Distant spread: Terms that highlight the extent of cancer spread beyond local regions.
  • Metachronous, Hematogenous spread, and Lymphatic spread: Specific pathways and timings of metastatic spread.
  • Micrometastases and Infiltration: Refer to tiny, often microscopic spread of cancer cells and their penetration into surrounding tissues.
  • Tumor spread and Extra-nodal extension: Describe the progression of the tumor and the extension beyond the lymph nodes.
Appendix D Prompt Used for Prediction
D.1 Zero-Shot Learning
Zero-shot learning task Task: Classify the presence of metastasis from patient clinical notes. Respond only with the following numerical codes based on the notes provided: - (1) Yes: The notes explicitly confirm the patient has metastasis. - (2) No: The notes explicitly confirm the patient does not have metastasis. - (3) Unknown: The notes do not contain sufficient information to determine the patient’s metastasis status. Instructions: 1. Do not provide explanations or reasons for your classification. 2. Use only the information in the notes for your classification. 3. If the notes are ambiguous or lack details regarding metastasis, choose "(3) Unknown". 4. Only provide a single numerical code as the response. Examples: - Example 1: "The CT scan shows multiple nodules in the liver consistent with metastasis." - (1) - Example 2: "Patient has a history of cancer but no evidence of metastatic disease on recent imaging." - (2) - Example 3: "Patient treated for localized breast absent of metastasis." - (3) Analyze the following clinical notes to determine if the patient has metastasis: {INSERT CURRENT CLINICAL NOTES HERE}
D.2 Few-Shot Learning
Few-shot learning task Task: Classify the presence of metastasis from patient clinical notes. Respond only with the following numerical codes based on the notes provided: - (1) Yes: The notes explicitly confirm the patient has metastasis. - (2) No: The notes explicitly confirm the patient does not have metastasis. - (3) Unknown: The notes do not contain sufficient information to determine the patient’s metastasis status. Instructions: 1. Do not provide explanations or reasons for your classification. 2. Use only the information in the notes for your classification. 3. If the notes are ambiguous or lack details regarding metastasis, choose "(3) Unknown". 4. Only provide a single numerical code as the response. Examples: {INSERT THE EXAMPLES HERE} Analyze the following clinical notes to determine if the patient has metastasis: {INSERT CURRENT CLINICAL NOTES HERE}